### 08 Jan 2024 v1

@@ 3 Learning and Knowledge in Cosmology
It is pointless to think of different universes with different laws of nature because they are beyond our comprehension. Therefore, following the reasonable little question of [16], we reduce the set of possible universes to those with the same laws but possibly different constants. In such scenario, there is a one-to-one relation between $\mathcal{X}$ and the multiverse because each $x\in\mathcal{X}$ represents a universe with value $x$ for the parameter $X$. In other words, there is a bijective function from $\mathcal{X}$ to the set of possible universes [4, 15, 48, 49, 50, 51, 52, 53, 54], with $x_{0}$ being the actual value of the constant of nature in our world, and any other $x\in\mathcal{X}$ being a counterfactual universe.
@@
can be fully learned and known. A sufficient condition for learning and knowledge acquisition of $p_{1}$ compared to an ignorant person is having $\mathrm{TP}_{\max}<1$.
@@
Identify the set of possible worlds $\Omega$ for the LKA problem with the set of possible universes $\mathcal{X}$, with $\omega_{0}$, the true world for the LKA problem, corresponding the observed value $x_{0}$ of $X$ for our universe. According to Section 2, the set $A=A_{1}$ of worlds for which $p_{1}$ is true coincides with $\ell_{X}$, the LPI of $X$.
@@
Then, according to Part II) of Definition 2 and Remark 2, we fully learn $p_{1}$ because ${\mathbf{P}}(A)=1$ and $p_{1}$ is true. Since ${\mathbf{P}}=\delta_{x_{0}}$, it follows from Part II) of Definition 3 that we also acquire full knowledge about $p_{1}$. This completes the proof of the first part of Theorem 1
For a proof of the second part of Theorem 1, we also need to consider the beliefs ${\mathbf{P}}_{0}$ of the ignorant person about the value of $X$. Recall that ${\mathbf{P}}_{0}$ is in maximum entropy over $\Omega=\mathcal{X}$, given the restrictions imposed by $\mathcal{F}$, the class of distributions considered in Table 1. This implies ${\mathbf{P}}_{0}\in\mathcal{F}$, and consequently
@@
After having dealt with learning and knowledge acquisition of our universe harboring life, let us now study whether it is possible to learn and know whether our universe is fine-tuned. This corresponds to the proposition
@@
of Algorithm 1 is true. The active information, for the set of worlds $A$ for which $p_{2}$ is true, then satisfies
$I^{+}(A)=\min(\log(\mathrm{TP}_{\mathrm{max}}^{-1},\delta^{-1}),$ (10)
@@
is enough to learn $p_{2}$ that $X$ is fine-tuned for life, whereas knowing $p_{2}$ requires more. For instance, (A) is a sufficient additional condition to assure that knowledge about $p_{2}$ has been acquired.
@@
Since $p_{2}$ is true by assumption, learning follows from Part I) of Definition 2 and the fact that $I^{+}(A)>0$, which follows from (10), (11) and $\delta<1$. Note that although the agent’s belief in (A) was used in order to construct the posterior belief ${\mathbf{P}}$ in Theorem 2, it is not required that (A) is true in order to derive ${\mathbf{P}}$ and formula (10) for $I^{+}(A)$.
If only (1) and (11), but not (A), hold, then it may happen that $\text{TP}_{\text{max}}<\omega_{0}\leq\delta$, so that for small enough $\epsilon>0$ we have $0={\mathbf{P}}(B_{\epsilon}[\omega_{0}])<2\epsilon={\mathbf{P}}_{0}(B_{\epsilon}[\omega_{0}])$. Whenever this happens, Condition 3 of Definition 3 is violated. On the other hand, if (1), (11), and (A) hold it is clear that $\omega_{0}\in(0,\text{TP}_{\text{max}})$. Since ${\mathbf{P}}_{0}\sim U(0,1)$ and ${\mathbf{P}}\sim U(0,\text{TP}_{\text{max}})$, it follows from Definition 3 that the agent has acquired knowledge about $p_{2}$ compared to the ignorant person. ∎
@@
Note that even though we assumed the agent’s belief in (A) in order to construct ${\mathbf{P}}$, this condition is not needed in order for him to learn about FT. Comparing Theorem 1 with Corollary 2.1, corresponding with intuition, we find that more is required to learn that our universe is fine-tuned ($p_{2}$) than to learn that it harbors life ($p_{1}$).
@@
are sufficient for establishing, learning and knowing that $X$ is fine-tuned for life.
@@
Since (A) and (B) imply (1), they are sufficient to establish FT (that $p_{2}$ is true). Also, (B) implies (11). Therefore, according to Corollary 2.1, it follows that the agent has learnt and acquired knowledge about FT compared to an ignorant person. ∎
@@
As seen in the proof of Corollary 2.1, it is possible to learn FT of $X$ for the wrong reasons if (1) and (B) hold but (A) does not. However, this is highly unsatisfactory because it does not tell us anything about the real probability $\omega_{0}=F_{0}(\ell_{X})$ of FT. Indeed, it is possible to falsely learn FT if (1) is false while (11) is true. Therefore, if FT cannot be known, it is not very useful. For this reason, we only pay attention to scenarios where FT can be known.
@@ 4.1 The crucial choice of $\mathcal{F}$
Theorem 3 says that it is enough to focus on (A) and (B), since these two equations imply that we can know that $X$ is fine-tuned for life. In order to analyse whether (A) and (B) are restrictive, recall that the first two steps of Algorithm 1 determine the size of the set $\cal F$ of prior distributions of $X$, which represent possible beliefs of an ignorant person about our universe. It turns out that these two steps are of crucial importance for the validity of (A) and (B), and hence for establishing that our universe is fine-tuned. The less we assume in Steps 1 and 2 of Algorithm 1 (the larger $\mathcal{F}$ is), the easier it is to satisfy (A), but the more difficult it is to satisfy (B). This has interesting implications for the list of possible choices of $\mathcal{F}$ in Table 1:
@@
On the other hand, Rows 3, 4, 6, and 7 of Table 1 represent a second scenario where $\mathcal{F}$ is smaller, making it easy to satisfy (B), but more difficult to satisfy (A). For instance, when the signal-to-noise ratio of the distribution $F_{X}$ of $X$ is bounded (Rows 3, 6, and 7), it is still possible that (A) does not hold (e.g., in the $\mathcal{X}=\mathbb{R}$ case, if the true maxent distribution $F_{0}$ of $X$ is positive, with a very small standard deviation in relation to its mean).
@@ 4.2 Random distributions
has low probability. (Note that $P(E)$ being small is less stringent than (A) failing with a small probability). Showing that $P(E)$ is small thus requires choosing the distribution of $F=F(\cdot;\theta)$, for which a distribution $G_{\theta}$ of $\theta$ is needed. This requires implementing Algorithm 1 on a new class of hyperhyperparameters $\lambda$ that parametrize the distribution $G_{\theta}=G(\cdot;\lambda)$ of $\theta$, which in turn leads to new restrictions $\lambda\in\Lambda$ on the hyperhyperparameter. We thus end up in an endless loop [55] because it requires showing that the hyperparameters $\theta$ themselves are tuned in order to obtain FT of the physical parameters $X$. In other words, we replace the FT question of the physical parameters $X$ with a tuning question of the hyperparameters $\theta$.
@@ 4.3 Bounded parameter space
This implies that for a finite parameter space $\cal X$ we can know that $X$ is fine-tuned, with a large probability $1-P(E)$, if $\cal F$ is chosen small enough so that (B) holds, but on the other large enough so that (13) is valid.
Since a finite parameter space is a favorite scenario for physicists (remember the discussion after Theorem 1), it is worth exploring whether there are actual FT instances in which the parameter space is actually finite. However, current arguments in favor of this hypothesis are unconvincing [11, 12]. Therefore, we need to deal with $\mathcal{X}$ being unbounded, and for such cases FT cannot be known unless some restrictions such as (A) and (B) are imposed.
@@ 4.4 The real line as parameter space
The argument of Section 4.4.3, based on random measures, is unfortunately not straightforward to generalize when the parameter space $\mathcal{X}$ of $X$ equals $\mathbb{R}$. Rows 4 and 5 of Table 1 correspond to a scale family of distributions that typically are symmetric around the origin. However, the scenario of Row 4 does not consider the observed value $x_{0}$ of the parameter in our universe as an estimator of a location parameter for $F_{X}$, making it plausible to suggest that (A) is violated. Indeed, it is a basic principle to obtain maxent distributions, subjecting them to the available observations [56, 57, 58, 59]. Thus, defining $\mathcal{F}$ without including into it a random variable whose location parameter is given by the observed value of $X$ in our universe might be too restrictive.
On the other hand, the location-scale family with an upper bound $T$ of the signal-to-noise ratio (Row 7 of Table 1), does include distributions with $x_{0}$ as mean parameter. It is possible then to know that $X$ is fine-tuned if the true distribution $F_{0}$ satisfies $\mbox{SNR}\leq T$ (so that (A) holds), and relative half length $\epsilon$ of the LPI is small enough so that (B) holds. The assumption $\mbox{SNR}\leq T$ cannot be derived from the maximum entropy principle though, and it rather needs some other justification (such as simplicity).
@@ 4.5 Positive line as parameter space
When $\mathcal{X}=\mathbb{R}^{+}$, it is possible to use Row 1 of Table 1. Notice that this is a scale family of distributions over $\mathbb{R}^{+}$ that includes all the exponential distributions. This is important because the exponential distribution $F(\cdot;\theta)$ is maxent over all distributions restricted to have mean $\theta$. Other maxent distributions over $\mathbb{R}^{+}$ are possible under different constraints; however, they are less parsimonious because they require more restrictions. For instance, the $\chi^{2}$ distribution of $X$ has maxent distribution over all distributions on the nonnegative reals given restrictions on $E(X)$ and $E(\log X)$ [see, e.g., 60], thus pertaining to Row 2 of Table 1. Hence, the family of exponential distributions $\{F(\cdot;\theta)\}_{\theta\in\mathbb{R}^{+}}$ is the less restrictive family of maximum entropy distributions over $\mathbb{R}^{+}$, and we can naturally include into it the exponential distribution with mean value corresponding to the observed value $x_{0}$ of the constant in our universe. Therefore, in this scenario it is plausible that (A) is satisfied and, whenever (B) holds as well, we can establish and know FT. In order to give a more explicit formulation of (B) for the family of exponential distributions of $X$ (cf. Row 1 of Table 1), we notice that the maximum tuning probability equals
@@ 5 Examples
We consider several examples of physical parameters $X$ to determine whether FT can be known for them. We will only treat cases in which the parameter cannot take negative values, and make use of Principle 1. This Section does not intend to be exhaustive. It rather intends to show two things. First, that FT can be known in a few instances. Second, that we can fail to know the level of tuning for a parameter if the relative half length $\epsilon$ of its LPI is not small. More bounds for the LPIs can be found for instance in [4] and references therein. In particular, Table 2 of [4] shows that many LPIs span over several orders of magnitude. This implies that FT cannot be known in those cases because $\epsilon$ is large and therefore (by Table 1) (B) is not satisfied.
@@ 5.1 Critical density of the universe
According to Paul [6], the critical density of the universe $\rho_{\text{crit}}$ cannot take values outside the interval
@@
where $\rho_{\text{crit,0}}$ is the observed critical density. Since the density cannot be negative, $\mathcal{X}=\mathbb{R}^{+}$. Taking $\epsilon=10^{-60}$, we notice that $\epsilon\ll 1$. According to row 1 of Table 1, [22] and (14), we have that $\text{TP}_{\text{max}}=2e^{-1}\times 10^{-60}$ is a small number. Thus, by Principle 1, we know that there is FT.
@@ 5.2 Gravitational force
According to Paul [6], when observing the ratio $X$ of the gravitational constant $G$ to the contribution from vacuum energy to the cosmological constant $\Lambda_{\text{vac}}$, gravitation cannot fall outside the life-permitting interval
@@
where $x_{0}$ is the observed value of $X$. Since gravitation is an attractive force, it cannot be negative, and therefore $X$ is positive as well (so that $\mathcal{X}=\mathbb{R}^{+}$). Since $\epsilon=10^{-100}$, it follows that $\epsilon\ll 1$. Thus, according to Row 1 of Table 1, [21] and (14), the maximal tuning probability is $\text{TP}_{\text{max}}=2e^{-1}\times 10^{-100}$. By Principle 1, we know that there is FT.
Notice that some physicists have considered the possibility that the gravitational force can also be repulsive [a possibility considered, for instance, in [5] and [1]]. In this scenario, additional assumptions beyond Principle 1 are required in order to know that $X$ is fine-tuned.

### 01 Mar 2024 v2

@@ 3 Learning and Knowledge in Cosmology
It is pointless to think of different universes with different laws of nature because they are beyond our comprehension. Therefore, following the reasonable little question of Barnes (2019-2020), we reduce the set of possible universes to those with the same laws but possibly different constants. In such scenario, there is a one-to-one relation between $\mathcal{X}$ and the multiverse because each $x\in\mathcal{X}$ represents a universe with value $x$ for the parameter $X$. In other words, there is a bijective function from $\mathcal{X}$ to the set of possible universes (Adams, 2019; Sandora, 2019a, b, c, d; Sandora et al., 2022, 2023a, 2023b, 2023c), with $x_{0}$ being the actual value of the constant of nature in our world, and any other $x\in\mathcal{X}$ being a counterfactual universe.
@@
can be fully learned and known. A sufficient condition for learning and knowledge acquisition of $p_{1}$, compared to an ignorant person, is having $\mathrm{TP}_{\max}<1$.
@@
Identify the set of possible worlds $\Omega$ for the learning and knowledge acquisition problem with the set of possible universes $\mathcal{X}$, with $\omega_{0}$ corresponding the observed value $x_{0}$ of $X$ for our universe. According to Section 2, the set $A=A_{1}$ of worlds for which $p_{1}$ is true coincides with $\ell_{X}$, the LPI of $X$.
@@
Then, according to Part II) of Definition 2 and Remark 2, we fully learn $p_{1}$ because ${\mathbf{P}}(A)=1$ and $p_{1}$ is true. Since ${\mathbf{P}}=\delta_{x_{0}}$, it follows from Part II) of Definition 3 that we also acquire full knowledge about $p_{1}$. This completes the proof of the first part of Theorem 1.
For a proof of the second part of Theorem 1, we also need to consider the beliefs ${\mathbf{P}}_{0}$ of the ignorant person about the value of $X$. Recall that ${\mathbf{P}}_{0}$ is in maxent over $\Omega=\mathcal{X}$, given the restrictions imposed by $\mathcal{F}$, the class of distributions considered in Table 1. This implies ${\mathbf{P}}_{0}\in\mathcal{F}$, and consequently
@@
After having dealt with learning and knowledge acquisition of our universe harboring life, let us now study whether it is possible to learn and know that our universe is fine-tuned. This corresponds to the proposition
@@
of Algorithm 1 is true. The active information for the set of worlds $A$ for which $p_{2}$ is true satisfies
$I^{+}(A)=\min(\log\mathrm{TP}_{\mathrm{max}}^{-1},\delta^{-1}),$ (10)
@@
is enough to learn $p_{2}$, that $X$ is fine-tuned for life, whereas knowing $p_{2}$ requires more. For instance, (A) is a sufficient additional condition to assure that knowledge about $p_{2}$ has been acquired.
@@
Since $p_{2}$ is true by assumption, learning follows from Part I) of Definition 2 and the fact that $I^{+}(A)>0$, which follows from (10), (11) and $\delta<1$. Note that although the agent’s belief in (A) was used to construct the posterior belief ${\mathbf{P}}$ in Theorem 2, it is not required that (A) is true to derive ${\mathbf{P}}$ and the formula in (10) for $I^{+}(A)$.
If only (1) and (11), but not (A), hold, then it may happen that $\text{TP}_{\text{max}}<\omega_{0}\leq\delta$, so that for small enough $\epsilon>0$ we have $0={\mathbf{P}}(B_{\epsilon}[\omega_{0}])<2\epsilon={\mathbf{P}}_{0}(B_{\epsilon}[\omega_{0}])$. Whenever this happens, Condition 3 of Definition 3 is violated. On the other hand, if (1), (11), and (A) hold, it is clear that $\omega_{0}\in(0,\text{TP}_{\text{max}})$. Since ${\mathbf{P}}_{0}\sim U(0,1)$ and ${\mathbf{P}}\sim U(0,\text{TP}_{\text{max}})$, it follows from Definition 3 that the agent has acquired knowledge about $p_{2}$ compared to the ignorant person. ∎
@@
Note that even though we assumed the agent’s belief in (A) to construct ${\mathbf{P}}$, this condition is not needed for him to learn about FT. Comparing Theorem 1 with Corollary 2.1, corresponding with intuition, we find that more is required to learn that our universe is fine-tuned ($p_{2}$) than to learn that it harbors life ($p_{1}$).
@@
are sufficient for establishing, learning, and knowing that $X$ is fine-tuned for life.
@@
Since (A) and (B) imply (1), they are sufficient to establish FT (that $p_{2}$ is true). Also, (B) implies (11). Therefore, according to Corollary 2.1, it follows that the agent has learnt and known FT compared to an ignorant person. ∎
@@
As seen in the proof of Corollary 2.1, it is possible to learn FT of $X$ for the wrong reasons if (1) and (B) hold but (A) does not. However, this is unsatisfactory because it does not say anything about the real probability $\omega_{0}=F_{0}(\ell_{X})$ of FT. Indeed, it is possible to falsely learn FT if (1) is false while (11) is true. Therefore, if FT cannot be known, learning is not useful. For this reason, we only pay attention to scenarios where FT can be known.
@@ 4.1 The crucial choice of $\mathcal{F}$
Theorem 3 says that it is enough to focus on (A) and (B), since these two equations imply that we can know that $X$ is fine-tuned for life. To analyse whether (A) and (B) are restrictive, recall that the first two steps of Algorithm 1 determine the size of the set $\cal F$ of prior distributions of $X$, which represent possible beliefs of an ignorant person about our universe. It turns out that these two steps are of crucial importance for the validity of (A) and (B), and hence for establishing that our universe is fine-tuned. The less we assume in Steps 1 and 2 of Algorithm 1 (the larger $\mathcal{F}$ is), the easier it is to satisfy (A), but the more difficult it is to satisfy (B). This has interesting implications for the list of possible choices of $\mathcal{F}$ in Table 1:
@@
On the other hand, Rows 3, 4, 6, and 7 of Table 1 represent a second scenario where $\mathcal{F}$ is smaller, making it easy to satisfy (B) but more difficult to satisfy (A). For instance, when the signal-to-noise ratio of the distribution $F_{X}$ of $X$ is bounded (Rows 3, 6, and 7), it is still possible that (A) does not hold (e.g., in the $\mathcal{X}=\mathbb{R}$ case, if the true maxent distribution $F_{0}$ of $X$ is positive, with a very small standard deviation in relation to its mean).
@@ 4.2 Random distributions
has low probability. (Note that $P(E)$ being small is less stringent than (A) failing with a small probability). Showing that $P(E)$ is small thus requires choosing the distribution of $F=F(\cdot;\theta)$, for which a distribution $G_{\theta}$ of $\theta$ is needed. This requires implementing Algorithm 1 on a new class of hyperhyperparameters $\lambda$ that parametrize the distribution $G_{\theta}=G(\cdot;\lambda)$ of $\theta$, which in turn leads to new restrictions $\lambda\in\Lambda$ on the hyperhyperparameter. We thus end up in an endless loop (Hofstadter, 1999) because it requires showing that the hyperparameters $\theta$ themselves are tuned to obtain FT of the physical parameters $X$. In other words, we replace the FT question of the physical parameters $X$ with a tuning question of the hyperparameters $\theta$.
@@ 4.3 Bounded parameter space
This implies that, for a finite parameter space $\cal X$, we can know that $X$ is fine-tuned, with a large probability $1-P(E)$, if $\cal F$ is chosen small enough so that (B) holds, but also large enough so that (13) is valid.
Since a finite parameter space is a favorite scenario for physicists (remember the discussion after Theorem 1), it is worth exploring whether there are actual FT instances in which the parameter space is actually finite. However, current arguments in favor of this hypothesis are unconvincing (Colyvan et al., 2005; McGrew et al., 2001). Therefore, we need to deal with $\mathcal{X}$ being unbounded, and for such cases FT cannot be known unless some restrictions, such as (A) and (B), are imposed.
@@ 4.4 The real line as parameter space
The argument of Section 4.4.3, based on random measures, is unfortunately not straightforward to generalize when the parameter space $\mathcal{X}$ of $X$ equals $\mathbb{R}$. Rows 4 and 5 of Table 1 correspond to a scale family of distributions that typically are symmetric around the origin. However, the scenario of Row 4 does not consider the observed value $x_{0}$ of the parameter in our universe as an estimator of a location parameter for $F_{X}$, making it plausible to suggest that (A) is violated. Indeed, it is a basic principle to obtain maxent distributions, subjecting them to the available observations (Celis et al., 2020; Singh & Vishnoi, 2014; Wainwright & Jordan, 2008; Suh, 2023). Thus, defining $\mathcal{F}$ without including into it a random variable whose location parameter is given by the observed value of $X$ in our universe might be too restrictive.
On the other hand, the location-scale family with an upper bound $T$ of the signal-to-noise ratio (Row 7 of Table 1) does include distributions with $x_{0}$ as mean parameter. It is possible then to know that $X$ is fine-tuned if the true distribution $F_{0}$ satisfies $\mbox{SNR}\leq T$ (so that (A) holds), and relative half length $\epsilon$ of the LPI is small enough so that (B) holds. The assumption $\mbox{SNR}\leq T$ cannot be derived from the maxent principle though, and it rather needs some other justification (such as simplicity).
@@ 4.5 Positive line as parameter space
When $\mathcal{X}=\mathbb{R}^{+}$, it is possible to use Row 1 of Table 1. Notice that this is a scale family of distributions over $\mathbb{R}^{+}$ that includes all the exponential distributions. This is important because the exponential distribution $F(\cdot;\theta)$ is maxent over all distributions restricted to have mean $\theta$. Other maxent distributions over $\mathbb{R}^{+}$ are possible under different constraints; however, they are less parsimonious because they require more restrictions. For instance, the $\chi^{2}$ distribution of $X$ has maxent distribution over all distributions on the nonnegative reals given restrictions on $E(X)$ and $E(\log X)$ (see, e.g., Park & Bera, 2009), thus pertaining to Row 2 of Table 1. Hence, the family of exponential distributions $\{F(\cdot;\theta)\}_{\theta\in\mathbb{R}^{+}}$ is the less restrictive family of maxent distributions over $\mathbb{R}^{+}$, and we can naturally include into it the exponential distribution with mean value corresponding to the observed value $x_{0}$ of the constant in our universe. Therefore, in this scenario it is plausible that (A) is satisfied and, whenever (B) holds as well, we can establish and know FT. To give a more explicit formulation of (B) for the family of exponential distributions of $X$ (cf. Row 1 of Table 1), we notice that the maximum tuning probability equals
@@ 5 Examples
We consider several examples of physical parameters $X$ to determine whether FT can be known for them. We will only treat cases in which the parameter cannot take negative values, and make use of Principle 1. This Section does not intend to be exhaustive. It rather intends to show two things. First, that FT can be known in a few instances. Second, that we can fail to know the level of tuning for a parameter if the relative half length $\epsilon$ of its LPI is not small. More bounds for the LPIs can be found for instance in (Adams, 2019) and references therein. In particular, Table 2 of (Adams, 2019) shows that many LPIs span over several orders of magnitude. This implies that FT cannot be known in those cases because $\epsilon$ is large and therefore (by Table 1) (B) is not satisfied.
@@ 5.1 Critical density of the universe
According to Davies (1982), the LPI of the critical density of the universe $\rho_{\text{crit}}$ is
@@
where $\rho_{\text{crit,0}}$ is the observed critical density. Since the density cannot be negative, $\mathcal{X}=\mathbb{R}^{+}$. Taking $\epsilon=10^{-60}$, we notice that $\epsilon\ll 1$. According to row 1 of Table 1, Díaz-Pachón et al. (2023) and (14), we have that $\text{TP}_{\text{max}}=2e^{-1}\times 10^{-60}$ is a small number. Thus, by Principle 1, we know that there is FT.
@@ 5.2 Gravitational force
According to Davies (1982), when observing the ratio $X$ of the gravitational constant $G$ to the contribution from vacuum energy to the cosmological constant $\Lambda_{\text{vac}}$, the LPI of the gravitational constant is
@@
where $x_{0}$ is the observed value of $X$. Since gravitation is attractive, it cannot be negative, and therefore $X$ is positive as well (so that $\mathcal{X}=\mathbb{R}^{+}$). Since $\epsilon=10^{-100}$, it follows that $\epsilon\ll 1$. Thus, according to Row 1 of Table 1, Díaz-Pachón et al. (2021) and (14), the maximal tuning probability is $\text{TP}_{\text{max}}=2e^{-1}\times 10^{-100}$. By Principle 1, we know that there is FT.